# hlinnerleftmargin=0, innerrightmargin=0, innertopmargin=0, innerbottommargin=0, backgroundcolor=lightyellow, linewidth=0 Ultra-Efficient Thermally Undercut Silicon Photonic Devices in a 300 mm CMOS Foundry Process

### v1

@@ Abstract
Silicon photonic devices fundamental to high-density wavelength-division multiplexed (DWDM) optical links and photonic switching networks, such as resonant modulators and Mach-Zehnder interferometers (MZIs), are highly sensitive to fabrication variations and operational temperature swings. However, thermal tuning to compensate for fabrication and operational temperature variations can result in prohibitive power consumption, challenging the scalability of energy-efficient photonic integrated circuits (PICs). In this work, we develop and demonstrate a wafer-scale thermal undercut in a 300 mm complementary metal oxide semiconductor (CMOS) foundry that dramatically improves the thermal isolation of thermo-optic devices by selectively removing substrate material beneath the waveguides and resonators. This approach significantly reduces the power required for thermal tuning across multiple device architectures, achieving almost a 5 improvement in tuning efficiency in a state-of-the-art 4.5 radius microdisk modulator and a 40 improvement in efficiency for a MZI phase shifter. Additionally, to the best of the authors’ knowledge, we demonstrate the first wafer-scale analysis of the impact of undercut trench geometry on device performance using comprehensive wafer-scale measurements across 64 reticles of a 300 mm silicon-on-insulator (SOI) wafer. These results highlight the effectiveness of this wafer-scale thermal undercut technique in minimizing energy consumption, paving the way for scalable, ultra low-power silicon photonic systems.
@@ Introduction
Resonant modulators, which are crucial for high-speed modulation in DWDM systems due to their inherent wavelength selectivity, benefit from reduced tuning power with the thermal undercut technique. We apply the wafer-scale thermal undercut to microring modulators, racetrack-style ring modulators, and state-of-the-art vertical junction microdisk modulators. Our experimental results demonstrate significant gains in tuning efficiency, achieving almost a 5 improvement for the microdisk modulators, while maintaining CMOS-compatible drive voltages. Linear thermo-optic phase shifters in MZIs, crucial for interleaving, switching, and routing in photonic networks, exhibit even greater reductions in power consumption, with over 40 improvement in tuning efficiency with optimized trench geometries. Furthermore, we demonstrate the capability of arbitrarily defined undercut openings, providing room for future improvement in all designs with optimized opening geometry.
To the best of the authors’ knowledge we present, for the first time, wafer-scale measurements of undercut microdisk modulators, illustrating tuning efficiency trends based on systematically varied undercut geometric parameters. By sweeping the dimensions and shapes of the undercut openings beneath these microdisk modulators, we observe that wider and more elongated trenches yield significantly improved thermal isolation and reduced power consumption. These measurements reveal the critical role of trench geometry in optimizing thermal performance, providing key insights for design strategies that maximize tuning efficiency across photonic systems. The reduction in power consumption achieved through this optimized undercut geometry has broad implications for DWDM links and other high-density photonic systems. In DWDM links, where minimizing power per channel is essential for scalability [29], the improved thermal efficiency allows for higher channel counts and lower energy-per-bit metrics, critical for data center and telecom applications. Similarly, in efficient large-scale photonic switching systems [30], the power savings gained through thermal undercut can alleviate the thermal load, enabling increased switching density with minimized power consumption. These findings establish wafer-scale thermal undercut as indispensable to advancing energy-efficient silicon photonics, addressing key challenges in power reduction for next-generation silicon photonic systems.
@@ Results
We design and demonstrate robust wafer-scale thermal undercut in collaboration with a 300 mm CMOS foundry [31], ensuring that the process can extend to high-volume manufacturing. As shown in Fig. 1a, design of experiment test structures with both circular and square trench geometries validate the ability to process a range of arbitrary shapes in a standard CMOS fabrication process. This flexibility is essential for optimizing thermal isolation across different device architectures and confirms the robustness of the undercut technique for varied applications. Comprehensive design rules ensure devices retain structural integrity when surrounded by undercut trenches. The thermal undercut process employed in this work involves selectively removing silicon substrate material [32]. This creates an air-filled trench that significantly enhances thermal isolation by lowering the effective thermal conductivity surrounding the active photonic device. Silicon, with a thermal conductivity of approximately 150 W/m·K[33, 34], provides a highly conductive pathway for heat dissipation, causing significant thermal leakage from the micro-heater to the bulk substrate. Further, the buried oxide has a thermal conductivity of approximately 1.4 W/m·K[35, 36], representing reduced thermal confinement. By introducing an undercut trench filled with air, which has a thermal conductivity of roughly 0.025 W/m·K[37], we reduce both heat dissipation pathways, effectively creating an insulating barrier around the device. This drastic reduction in thermal conductivity between the heater and the substrate allows the device to reach higher temperatures with lower input electrical power, significantly improving thermal tuning efficiency.
@@
Thermal simulations [38] were performed on a microdisk modulator both with and without undercut to compare their transient thermal responses [39]. A schematic representation of a microdisk modulator surrounded by thermal undercut is shown in Fig. 2b. The comprehensive device design is detailed in ref. [40], and preliminary results of the undercut device are shown in ref. [41]. Simulations in which 1 mW of power was applied through the integrated micro-heater show a significantly higher temperature increase in the undercut microdisk compared to the nominal device, underscoring the enhanced thermal isolation achieved by the trench (Fig. 2c). This heightened temperature response for the same input power demonstrates the potential for lower power consumption, as less energy is required to reach the desired tuning temperatures in devices with thermal undercut. Additionally, transient thermal simulations were performed to provide insight into the dynamic response of the undercut device, revealing longer rise and fall times compared to the nominal device. Specifically, the nominal device had a rise time of 11.0 s and a fall time of 11.3 s, while introducing the undercut resulted in a rise time of 96.7 s and a fall time of 95.7 s. The approximately 8-9 slower heating and cooling transients are a result of the reduced thermal conductivity around the microdisk due to the air gap created by the undercut. This effect can be advantageous in applications where slower thermal responses improve stability by reducing sensitivity to temperature fluctuations.
@@
Microdisk modulator test structures were measured across two 300 mm wafers. Two variants of the device were measured on each reticle: modulators both with and without thermal undercut. A micrograph of a microdisk modulator with thermal undercut is shown in Fig. 2a, which confirms that the device is fully released from the substrate. Further demonstrating the full release of the device, a micrograph of the side profile is shown in Fig. 2b. A dicing saw was used to make a precise cut close to the undercut modulator on a singulated die to enable capture of the cross-section. The optical spectrum of a nominal microdisk modulator without thermal undercut at different applied thermo-optic phase shifter powers showing the shift in resonance is displayed in Fig. 2c. A drastic increase in the resonance shift for similar applied powers is shown in Fig. 2d, where the optical spectrum of a representative thermally undercut microdisk modulator is displayed. This large increase in tuning efficiency is additionally observed at the wafer-scale; the wafer maps are shown in Fig. 2g-h, showing the of the nominal non-undercut and undercut microdisks, respectively. The for the undercut microdisks is on average 5.32 mW, which is an improvement of 4.92 as compared to the of 26.14 mW exhibited by the devices without undercut. Furthermore, these wafer-scale measurements confirm the structural integrity and high yield of the undercut devices.
@@
To demonstrate the wide applicability of the undercut process, two other types of resonant modulators were fabricated, both with and without thermal undercut. The first device is a lateral junction microring modulator with a radius of 20 . The micrograph of this microring modulator with thermal undercut is shown in Fig. 3a. The thermo-optic phase shifter is located concentrically inside the ring. Trench openings were placed both inside and outside the microring and alongside the access waveguide. The second device is a racetrack-style ring modulator; a micrograph of a representative device is shown in Fig. 3d. A linear thermo-optic phase shifter is located internal to the ring along the straight access section. The PN-junction for high-speed modulation is located along the opposite straight section. For this modulator design, trench openings were placed both internally and externally along the straight sections of the ring and alongside the access waveguide. The measured across a wafer is shown for the non-undercut microring modulator and racetrack ring modulator in Fig. 3b and Fig. 3e, respectively. The wafer maps of the undercut microring and racetrack modulators are displayed in Fig. 3c. and Fig. 3f, respectively. We see an average tuning efficiency improvement of 25.1, from 100.01 mW without undercut to 3.98 mW with undercut for the microring modulator. In the case of the racetrack-style ring modulator, the tuning efficiency improvement on average is 48.1, from 63.11 mW without undercut to 1.31 mW with undercut. Both undercut devices exhibit drastic increases in tuning efficiency, indicating full release from the substrate.
@@
While all the fully-released undercut resonant devices show substantial improvements in tuning efficiency, the improvement factor is more pronounced for less efficient thermo-optic phase shifter designs. The baseline tuning efficiencies of the microring and racetrack modulators without undercut have values of 100.01 mW and 63.11 mW, respectively. In comparison, the already relatively efficient microdisk modulator exhibits a of 26.14 mW without undercut. The higher baseline efficiency of the microdisk can be attributed to its integrated thermo-optic phase shifter, which is surrounded by only a thin 100 nm oxide layer for electrical isolation [42]. The proximity of the phase shifter to the optical mode, with silicon as the dominant material between them, ensures effective thermal conduction due to silicon’s higher thermal conductivity compared to oxide. In contrast, the thermo-optic phase shifters in the microring and racetrack modulators are positioned farther from the optical mode, separated by a larger oxide gap. This separation significantly reduces thermal efficiency and confinement in these devices compared to the microdisk. As a result, the microring and racetrack modulators benefit more, relatively speaking, from the enhanced thermal isolation provided by the substrate undercut, achieving greater improvement factors in tuning efficiency.
@@
In addition to resonant devices, we designed and fabricated linear thermo-optic phase shifters both with and without undercut. A micrograph of the phase shifter with undercut is shown in Fig. 4a. The phase shifter consists of two doped silicon strips parallel to the waveguide. To extract the tuning efficiency of the phase shifter, it was placed within an imbalanced MZI so the interference fringes could be tracked at varying heater powers. The transmission spectrum of the MZI without undercut and with undercut under different applied heater powers is shown in Fig. 4c and Fig. 4d, respectively. The undercut phase shifter requires significantly less power to induce a larger shift in interference fringe wavelength. This is demonstrated in the wafer maps showing for the phase shifters without undercut and phase shifters with undercut in Fig. 4g and Fig. 4h, respectively. An improvement factor of 40.5 is shown, moving from a non-undercut of 43.46 mW to a of 1.07 mW for the undercut device. To characterize the time-dependent transient effects of substrate undercut on thermal modulation, we apply a square wave to the phase shifter. The transient is shown for the phase shifter without undercut in Fig. 4e, and similarly Fig. 4f shows the transient for the phase shifter with undercut. A pulse width of 100 was applied to the non-undercut device, resulting in a rise time of 12.6 and a fall time of 25.3 . For the undercut device, a pulse width of 10 ms was used, resulting in a rise time of 1.853 ms and a fall time of 1.258 ms. The increase in rise/fall time has implications on the performance of these phase shifters in the context of optical switches and multiplexers. For instance, the increase in rise/fall time results in a slower optical switching speed. However, the reduced switching speed disadvantage is mostly negated if the network topology is configured before application runtime [43, 44]. Further, the increase in rise/fall time can be leveraged when considering time-multiplexed driving signals for many thermo-optic phase shifters to reduce electrical I/O [45].
@@
To elucidate the impact of undercut geometry on the tuning efficiency of devices, we swept the undercut geometric parameters across a number of resonant devices. A micrograph of the test structure sweep of undercut microdisk modulators is shown in Fig. 5a. While keeping the microdisk modulator design constant for all cases, both the undercut opening size and separation were varied according to the schematic in Fig. 5b. The tuning efficiency of each microdisk modulator in the undercut parameter sweep was extracted across two full wafers. The tuning efficiency of each modulator in the parameter sweep, averaged across the first wafer, is shown in Figs. 5c and 5d, where the top trench length is 20 and 40 , respectively. Similarly, the tuning efficiencies of the modulators from the second full wafer are shown in Figs. 5e and 5f. First, there is a clear trend in increased tuning efficiency (reduced ) in all cases as the trench width is increased from 10 to 25 . The larger trench opening increases the lateral undercut, ensuring the devices are fully released from the substrate. However, we see diminishing returns when increasing the trench width from 20 to 25 , suggesting the tuning efficiency of the devices is dominated by the volume of remaining buried oxide and substrate contacting the bottom of the device. Additionally, there is a trend in increased when the trench separation is increased, which corresponds to wider oxide bridges between the trench openings. This trend explains the limited gains in tuning efficiency seen in processes using a sealed substrate undercut [26, 46, 27]; heat from the phase shifter is less confined to the device without the trench openings. Another notable trend is the marked consistency in tuning efficiency of the undercut modulators between both wafers, demonstrating the robustness and reliability of the substrate removal process.
@@
The small size of microdisk modulators causes strong light confinement, leading to high optical power density and nonlinear effects like thermal bistability [47]. Thermal isolation exacerbates this by trapping heat in the microdisk, complicating high-power management in resonant modulators. We studied the effects of high optical power in thermally undercut microdisk modulators, and thermal instability even at -14 dBm input power is observed [48]. The modulator is thermally isolated using wafer-scale undercut, shown in Fig. 6a. We examined electro-optic modulation of the PN-junction at varying laser powers. At low input laser power (-17.5 dBm), the DC spectral response showed a modulation efficiency of 95 pm/V under bias voltages from 0.5 V to -2.5 V (Fig. 6b). However, increasing laser power caused nonlinear spectral shifts with resonance red-shifting and broadening due to thermal instability, as shown in Fig. 6c. Near-resonance laser light is partially absorbed via two-photon and free-carrier absorption, generating heat. This effect is enhanced by the thermal isolation of the undercut structure, leading to significant thermal instability at laser optical powers as low as -14 dBm (Fig. 6d).
@@
Despite this, robust high-speed modulation was achievable over a broad spectral range at higher powers, up to -4 dBm. Thermally-induced resonance red-shifting enabled modulation over a range 6 wider than the resonator’s passive linewidth. We present an experimental investigation of optical modulation in thermally undercut microdisk modulators, focusing on the advantages of optically induced thermal nonlinearity. As laser power increases, the spectral shift of the resonator becomes nonlinear, accompanied by a redshift and resonance broadening due to thermal instability. We also explored the modulation behavior across varying operating frequencies and input optical powers. A 2 V peak-to-peak square wave was applied at different frequencies, and the peak-to-peak optical modulation was measured and normalized to compare modulation depths. Across all modulation frequencies, the spectral range broadens with increased laser power. However, at low frequencies (<1 MHz, Fig. 6e), although the spectral range expands, the peak-to-peak modulation decreases at higher laser powers, likely due to the device’s thermal response. In contrast, at higher data frequencies (>25 MHz, Fig. 6f-h), the modulation depth remains stable as laser power increases, achieving modulation over a spectral range six times broader than the resonator’s linewidth at an optical power of -4 dBm. Therefore, at high data rates, the device’s thermal sensitivity becomes a beneficial nonlinear effect, enabling robust modulation across a wide spectral range. Thermal isolation thus enhances modulation efficiency and reduces micro-heater tuning requirements.

### v2

# hlinnerleftmargin=0, innerrightmargin=0, innertopmargin=0, innerbottommargin=0, backgroundcolor=lightyellow, linewidth=0 Highly Uniform Thermally Undercut Silicon Photonic Devices in a 300 mm CMOS Foundry Process

@@ Abstract
Silicon photonic devices fundamental to high-density wavelength-division multiplexed (DWDM) optical links and photonic switching networks, such as resonant modulators and Mach-Zehnder interferometers (MZIs), are highly sensitive to fabrication variations and operational temperature swings. However, thermal tuning to compensate for fabrication and operational temperature variations can result in prohibitive power consumption, challenging the scalability of energy-efficient photonic integrated circuits (PICs). In this work, we develop and demonstrate a wafer-scale thermal undercut process in a 300 mm complementary metal oxide semiconductor (CMOS) foundry that dramatically improves the thermal isolation of thermo-optic devices by selectively removing substrate material beneath the waveguides and resonators. This approach significantly reduces the power required for thermal tuning across multiple device architectures, achieving almost a 5 improvement in tuning efficiency in a state-of-the-art 4.5 radius microdisk modulator and a 40 improvement in efficiency for a MZI phase shifter. To the best of the authors’ knowledge, we demonstrate the first wafer-scale comparison of non-undercut and undercut silicon photonic devices using comprehensive wafer-scale measurements across 64 reticles of a 300 mm silicon-on-insulator (SOI) wafer. Further, we demonstrate a comprehensive wafer-scale analysis of the influence of undercut trench opening geometry on device tuning efficiency. Notably, we observe highly uniform performance across the full 300 mm wafer for multiple device types, emphasizing that our process can be scaled to large-scale photonic circuits with high yield. These results open new opportunities for large-scale integrated photonic circuits using thermo-optic devices, paving the way for scalable, low-power silicon photonic systems.
@@ Introduction
Resonant modulators, which are crucial for high-speed modulation in DWDM systems due to their inherent wavelength selectivity, benefit from reduced tuning power with the thermal undercut technique. We apply the wafer-scale thermal undercut to microring modulators, racetrack-style ring modulators, and state-of-the-art vertical junction microdisk modulators. Our experimental results demonstrate significant gains in tuning efficiency, achieving almost a 5 improvement for the microdisk modulators, while maintaining CMOS-compatible drive voltages. Linear thermo-optic phase shifters in MZIs, crucial for interleaving, switching, and routing in photonic networks, exhibit even greater reductions in power consumption, with over 40 improvement in tuning efficiency with optimized trench geometries. We show large-scale uniformity in performance for each device type, exhibiting the robustness of the substrate removal process across full 300 mm wafers. Furthermore, we demonstrate the capability of arbitrarily defined undercut openings, providing room for future improvement in all designs with optimized opening geometry.
To the best of the authors’ knowledge we present, for the first time, comprehensive wafer-scale comparisons of non-undercut and undercut silicon photonic devices across multiple device types and heater geometries. Additionally, we illustrate tuning efficiency trends of undercut microdisk modulators based on systematically varied undercut geometric parameters. By sweeping the dimensions and shapes of the undercut openings beneath these microdisk modulators, we observe that wider and more elongated trenches yield significantly improved thermal isolation and reduced power consumption. These measurements reveal the critical role of trench geometry in optimizing thermal performance, providing key insights for design strategies that maximize tuning efficiency across photonic systems. The reduction in power consumption achieved through this optimized undercut geometry has broad implications for DWDM links and other high-density photonic systems. In DWDM links, where minimizing power per channel is essential for scalability [29], the improved thermal efficiency allows for higher channel counts and lower energy-per-bit metrics, critical for data center and telecom applications. Similarly, in efficient large-scale photonic switching systems [30], the power savings gained through thermal undercut can alleviate the thermal load, enabling increased switching density with minimized power consumption.
@@ Results
We design and demonstrate robust wafer-scale thermal undercut in collaboration with a 300 mm CMOS foundry [31], ensuring that the process can extend to high-volume manufacturing. This thermal undercut process was designed to maintain full compatibility with the standard AIM Photonics Actives Process, as highlighted by the process cross section in Fig. 1a. As shown in Fig. 1b, design of experiment test structures with both circular and square trench geometries validate the ability to process a range of arbitrary shapes in a standard CMOS fabrication process. This flexibility is essential for optimizing thermal isolation across different device architectures and confirms the robustness of the undercut technique for varied applications. Comprehensive design rules ensure devices retain structural integrity when surrounded by undercut trenches. The thermal undercut process employed in this work involves selectively removing silicon substrate material [32]. This creates an air-filled trench that significantly enhances thermal isolation by lowering the effective thermal conductivity surrounding the active photonic device. Silicon, with a thermal conductivity of approximately 150 W/m·K[33, 34], provides a highly conductive pathway for heat dissipation, causing significant thermal leakage from the micro-heater to the bulk substrate. Further, the buried oxide has a thermal conductivity of approximately 1.4 W/m·K[35, 36], representing reduced thermal confinement. By introducing an undercut trench filled with air, which has a thermal conductivity of roughly 0.025 W/m·K[37], we reduce both heat dissipation pathways, effectively creating an insulating barrier around the device. This drastic reduction in thermal conductivity between the heater and the substrate allows the device to reach higher temperatures with lower input electrical power, significantly improving thermal tuning efficiency. Since the foundry process does not include a metal heater layer such as titanium nitride, we implement our resistive heaters using doped silicon in the silicon device layer (Fig. 1a). We use a suite of standard n-type and p-type ion-implanted doping layers [31] to modulate the resistance of the silicon microheater and form ohmic contacts between the metal vias and silicon. Depending on the heater geometry (length, width, etc.) and the target resistance for the heater, we employ a mix of low, medium, and high doping levels ( cm-3, cm-3, and cm-3, respectively). The metal vias and routing layers are fabricated using standard copper and aluminum back end of line (BEOL) processing [31].
@@
Thermal simulations [38] were performed for a microdisk modulator both with and without undercut to compare their transient thermal responses [39]. The mask layout for a microdisk modulator surrounded by thermal undercut is shown in Fig. 2a. The comprehensive device design is detailed in ref. [40], and preliminary results of the undercut device are shown in ref. [41]. Although outside the scope of this work, recent improvements to the microdisk modulator RF contact geometry have yielded devices capable of 32 Gb/s on-off-keying non-return-to-zero (OOK-NRZ) transmission with low peak-to-peak driving voltages [3]. Simulations in which 1 mW of power was applied through the integrated micro-heater show a significantly higher temperature increase in the undercut microdisk compared to the nominal device, underscoring the enhanced thermal isolation achieved by the trench (Fig. 2d). This heightened temperature response for the same input power demonstrates the potential for lower power consumption, as less energy is required to reach the desired tuning temperatures in devices with thermal undercut. Additionally, transient thermal simulations were performed to provide insight into the dynamic response of the undercut device, revealing longer rise and fall times compared to the nominal device. Specifically, the nominal device had a rise time of 11.0 s and a fall time of 11.3 s, while introducing the undercut resulted in a rise time of 96.7 s and a fall time of 95.7 s. The approximately 8-9 slower heating and cooling transients are a result of the reduced thermal conductivity around the microdisk due to the air gap created by the undercut. This effect can be advantageous in applications where slower thermal responses improve stability by reducing sensitivity to temperature fluctuations.
@@
Microdisk modulator test structures were measured across two 300 mm wafers. Two variants of the device were measured on each reticle: modulators both with and without thermal undercut. A micrograph of a microdisk modulator with thermal undercut is shown in Fig. 2c, which confirms that the device is fully released from the substrate. Further demonstrating the full release of the device, a micrograph of the side profile is shown in Fig. 1c. A dicing saw was used to make a precise cut close to the undercut modulator on a singulated die to enable capture of the cross-section. The optical spectrum of a nominal microdisk modulator without thermal undercut at different applied thermo-optic phase shifter powers showing the shift in resonance is displayed in Fig. 2e. A drastic increase in the resonance shift for similar applied powers is shown in Fig. 2f, where the optical spectrum of a representative thermally undercut microdisk modulator is displayed. This large increase in tuning efficiency is additionally observed at the wafer-scale; the wafer maps are shown in Figs. 2e and 2f, showing the of the nominal non-undercut and undercut microdisks, respectively. The for the undercut microdisks is on average 5.316 mW with standard deviation 0.046 mW, which is an improvement of 4.92 as compared to the of 26.136 mW with standard deviation 0.219 mW exhibited by the devices without undercut. Furthermore, these wafer-scale measurements confirm the structural integrity and high yield of the undercut devices.
@@
To demonstrate the wide applicability of the undercut process, two other types of resonant modulators were fabricated, both with and without thermal undercut. The first device is a lateral junction microring modulator with a radius of 20 . The micrograph of this microring modulator with thermal undercut is shown in Fig. 3b. The thermo-optic phase shifter is located concentrically inside the ring. Trench openings were placed both inside and outside the microring and alongside the access waveguide, as shown in the schematic in Fig. 3a. The second device is a racetrack-style ring modulator; a micrograph of a representative device is shown in Fig. 3f. A linear thermo-optic phase shifter is located internal to the ring along the straight access section. The PN-junction for high-speed modulation is located along the opposite straight section. For this modulator design, trench openings were placed both internally and externally along the straight sections of the ring and alongside the access waveguide, as shown in the schematic in Fig. 3e. The measured across a wafer is shown for the non-undercut microring modulator and racetrack ring modulator in Fig. 3c and Fig. 3g, respectively. The wafer maps of the undercut microring and racetrack modulators are displayed in Fig. 3d and Fig. 3h, respectively. We observe an average tuning efficiency improvement of 25.1, from average of 100.01 mW and standard deviation 4.46 mW without undercut to average of 3.99 mW and standard deviation 0.05 mW with undercut for the microring modulator. In the case of the racetrack-style ring modulator, the tuning efficiency improvement on average is 48.1, from average of 63.11 mW and standard deviation 3.43 mW without undercut to average of 1.31 mW and standard deviation 0.01 mW with undercut. Both undercut devices exhibit drastic increases in tuning efficiency, indicating full release from the substrate.
@@
While all the fully-released undercut resonant devices show substantial improvements in tuning efficiency, the improvement factor is more pronounced for less efficient thermo-optic phase shifter designs. The baseline tuning efficiencies of the microring and racetrack modulators without undercut have values of 100.01 mW and 63.11 mW, respectively. In comparison, the already relatively efficient microdisk modulator exhibits a of 26.136 mW without undercut. The higher baseline efficiency of the microdisk can be attributed to its integrated thermo-optic phase shifter, which is surrounded by only a thin 100 nm oxide layer for electrical isolation [42]. The proximity of the phase shifter to the optical mode, with silicon as the dominant material between them, ensures effective thermal conduction due to silicon’s higher thermal conductivity compared to oxide. In contrast, the thermo-optic phase shifters in the microring and racetrack modulators are positioned farther from the optical mode, separated by a larger oxide gap. This separation significantly reduces thermal efficiency and confinement in these devices compared to the microdisk. As a result, the microring and racetrack modulators benefit more, relatively speaking, from the enhanced thermal isolation provided by the substrate undercut, achieving greater improvement factors in tuning efficiency. Consequently, more conservative phase shifter designs, such as those designed to incur negligible optical loss, can still achieve state-of-the-art tuning efficiency once undercut.
@@
In addition to resonant devices, we designed and fabricated linear thermo-optic phase shifters both with and without undercut. The mask layout and a representative micrograph of the phase shifter with undercut are shown in Figs. 4a and 4b and Fig. 4c, respectively. The phase shifter consists of two doped silicon strips parallel to the waveguide. These doped silicon strips are placed at a distance away from the optical mode to avoid excess optical losses, resulting in the insertion loss being dominated by the intrinsic waveguide propagation loss. A 125 device length and 2 dB/cm propagation loss yields a total insertion loss of 0.025 dB. To extract the tuning efficiency of the phase shifter, it was placed within an imbalanced MZI so the interference fringes could be tracked at varying heater powers. The transmission spectrum of the MZI without undercut and with undercut under different applied heater powers is shown in Fig. 4e and Fig. 4f, respectively. The undercut phase shifter requires significantly less power to induce a larger shift in interference fringe wavelength. This is demonstrated in the wafer maps showing for the phase shifters without undercut and phase shifters with undercut in Fig. 4e and Fig. 4f, respectively. An improvement factor of 40.5 is shown, moving from a non-undercut average of 43.46 mW and standard deviation 0.94 mW to an average of 1.07 mW and standard deviation 0.01 mW for the undercut device. To characterize the time-dependent transient effects of substrate undercut on thermal modulation, we apply a square wave to the phase shifter. The transient is shown for both the nominal and undercut phase shifter in Fig. 4d (top and bottom, respectively). A pulse width of 100 was applied to the non-undercut device, resulting in a rise time of 12.6 and a fall time of 25.3 . For the undercut device, a pulse width of 10 ms was used, resulting in a rise time of 1.853 ms and a fall time of 1.258 ms. The increase in rise/fall time has implications on the performance of these phase shifters in the context of optical switches and multiplexers. For instance, the increase in rise/fall time results in a slower optical switching speed. However, the reduced switching speed disadvantage is mostly negated if the network topology is configured before application runtime [43, 44]. Further, the increase in rise/fall time can be leveraged when considering time-multiplexed driving signals for many thermo-optic phase shifters to reduce electrical I/O [45].
@@
To elucidate the impact of undercut geometry on the tuning efficiency of devices, we swept the undercut geometric parameters across a number of microdisk modulator devices. A micrograph of the test structure sweep of undercut microdisk modulators is shown in Fig. 5a. While keeping the microdisk modulator design constant for all cases, both the undercut opening size and separation were varied according to the schematic in Fig. 5b. The tuning efficiency of each microdisk modulator in the undercut parameter sweep was extracted across two full wafers. The tuning efficiency of each modulator in the parameter sweep, averaged across the first wafer, is shown in Figs. 5c and 5d, where the top trench length is 20 and 40 , respectively. Similarly, the tuning efficiencies of the modulators from the second full wafer are shown in Figs. 5e and 5f. First, there is a clear trend in increased tuning efficiency (reduced ) in all cases as the trench width is increased from 10 to 25 . The larger trench opening increases the lateral undercut, ensuring the devices are fully released from the substrate. However, we see diminishing returns when increasing the trench width from 20 to 25 , suggesting the tuning efficiency of the devices is dominated by the volume of remaining buried oxide and substrate contacting the bottom of the device. Additionally, there is a trend in increased when the trench separation is increased, which corresponds to wider oxide bridges between the trench openings. This trend explains the limited gains in tuning efficiency seen in processes using a sealed substrate undercut [26, 46, 27]; heat from the phase shifter is less confined to the device without the trench openings. Another notable trend is the marked consistency and uniformity in tuning efficiency of the undercut modulators between both wafers, demonstrating the mechanical robustness and reliability of the substrate removal process. Only when aggressively pushing the boundaries of trench separation far beyond the foundry design rules do we see mechanical damage to the bus waveguide and metal traces. Using Fig. 5b as a reference, we observed mechanical damage at a trench separation (TS) of 1 , but observed no damage and high mechanical stability at a trench separation of at least 2 .
@@
The small size of microdisk modulators causes strong light confinement, leading to high optical power density and nonlinear effects like thermal bistability [47]. Thermal isolation exacerbates this by trapping heat in the microdisk, complicating high-power management in resonant modulators [48]. We studied the effects of high optical power in thermally undercut microdisk modulators, and thermal instability even at -14 dBm input power is observed [49]. The modulator is thermally isolated using wafer-scale undercut, shown in Fig. 6a. We examined electro-optic modulation of the PN-junction at varying laser powers. At low input laser power (-17.5 dBm), the DC spectral response showed a modulation efficiency of 65 pm/V under bias voltages from 0.5 V to -2.5 V (Fig. 6b). However, increasing laser power caused nonlinear spectral shifts with resonance red-shifting and broadening due to thermal instability, as shown in Fig. 6c [50]. Near-resonance laser light is partially absorbed via two-photon and free-carrier absorption, generating heat. This effect is enhanced by the thermal isolation of the undercut structure, leading to significant thermal instability at laser optical powers as low as -14 dBm (Fig. 6d).
@@
Despite this, robust high-speed modulation was achievable over a broad spectral range at higher powers, up to -4 dBm. Thermally-induced resonance red-shifting enabled modulation over a range 6 wider than the resonator’s passive linewidth. We present an experimental investigation of optical modulation in thermally undercut microdisk modulators, focusing on the advantages of optically induced thermal nonlinearity. As laser power increases, the spectral shift of the resonator becomes nonlinear, accompanied by a redshift and resonance broadening due to thermal instability. We also explored the modulation behavior across varying operating frequencies and input optical powers. A 2 V peak-to-peak square wave was applied at different frequencies, and the peak-to-peak optical modulation was measured and normalized to compare modulation depths. Across all modulation frequencies, the spectral range broadens with increased laser power. However, at low frequencies (<1 MHz, Fig. 6e), although the spectral range expands, the peak-to-peak modulation decreases at higher laser powers, likely due to the device’s thermal response. In contrast, at higher data frequencies (>25 MHz, Figs. 6f-h), the modulation depth remains stable as laser power increases, achieving modulation over a spectral range six times broader than the resonator’s linewidth at an optical power of -4 dBm. Therefore, at high data rates, the device’s thermal sensitivity becomes a beneficial nonlinear effect, enabling robust modulation across a wide spectral range. Thermal isolation thus enhances modulation efficiency and reduces micro-heater tuning requirements.